Structuring a Make-or-Buy Problem Fresh Foods, a large restaurant chain, needs to determine if it would...

  1. Structuring a Make-or-Buy Problem

    Fresh Foods, a large restaurant chain, needs to determine if it would be cheaper to produce 5,000 units of its main food ingredient for use in its restaurants or to purchase them from an outside supplier for $12 each. Cost information on internal production includes the following:

    Total Cost Unit Cost
    Direct materials $25,000    $5.00   
    Direct labor 15,000    3.00   
    Variable manufacturing overhead 7,500    1.50   
    Variable marketing overhead 11,500    2.30   
    Fixed plant overhead 30,000    6.00   
    Total $89,000    $17.80   

    Fixed overhead will continue whether the ingredient is produced internally or externally. No additional costs of purchasing will be incurred beyond the purchase price. If required, round your answers to the nearest whole number.

    Required:

    1. What are the alternatives for Fresh Foods?
    Make the ingredient in house or buy it externally.

    2. List the relevant cost(s) of internal production and of external purchase.

    All of the above

    3. Which alternative is more cost effective and by how much? (Use total cost when giving your answer.)

    Make $

    4. Now assume that 20% of the fixed overhead can be avoided if the ingredient is purchased externally. Which alternative is more cost effective and by how much? (Use total cost when giving your answer.)

    Buy $
Answer #1

Alternatives are:

Make the ingredient in house or buy it externally.

2.

Relevant costs are as follows:

Make

Buy

Direct material

5

Direct labor

3

Variable manufacturing overhead

1.5

Variable marketing overhead

2.3

Purchase cost

12

Relevant costs

11.8

12

Fixed overheads are not relevant since they will be incurred under both the alternatives.

3.Make is effective by (12-11.8)*5000 = $1,000

4.Cost of making now = 11.8 + 6*20% = $13

Buy is effective by $5000